Please I need answer for question number 3
I need the answer typed not handwriting please

T-Mobile 令 8:55 PM a moodle.mcny.edu Consider the supply of computers. For each of the a. A change in technology that lowers production 2. following, state the effect on supply: costs b. An increase in the price of semiconductors c. A decrease in the price of computers d. An increase in the wages of computer assembly workers e. An increase in consumer incomes 3. The demand curve is given by Qp500 50.5 10Py- 2Pz where p quantity demanded of good X Py price of good X I- consumer income, in thousands Py = price of good Y a. Based on the demand curve above, is X a nor- b. Based on the demand curve above, what is the c. Based on the demand curve above, what is the d. What is the equation of the demand curve if Pz- price of good Z mal or an inferior good? relationship between good X and good Y? relationship between good X and good Z? consumer incomes are $30,000, the price of e. Graph the demand curve that you found in (d), f. If the price of good Xis $15, what is the quantity g. Now suppose the price of good Y rises to $15. good Y is $10, and the price of good Z is $20? showing intercepts and slope demanded? Show this point on your demand curve Graph the new demand curve
Answer #1

3. a) Normal good is the good whose income effect is positive. Increase in income increases demand of good and vice-versa.

Inferior good is *the good which has negative income effect. Increase in income decreases demand of good and vice-versa.

Increase in I causes increase in QD so good is Normal Good.

b) Increase in price of good Y causes increase in quantity demanded of good X. Because dQd/dPy = 10 i.e. positive

There is direct relationship between good X and Y. Increase in price of Y increases demand of good X.

c) dQd/dPz = - 2 i.e. negative

There is inverse relationship between good X and Z. Increase in price of Z decreases demand of good X.

d) Qd = 500 - 5Px + 0.5(30000) + 10(10) - 2(20)

Qd = 500 - 5Px + 15000 + 100 - 40

Qd = 15560 - 5Px

e) When Px = 0 then Qd = 15560

When Qd = 0 then Px = 15560/5 = 3112

Slope of demand curve; dQd/dPx = - 5